# GenRec: Unifying Video Generation and Recognition with Diffusion Models

### v1

@@ Abstract
Video diffusion models are able to generate high-quality videos by learning strong spatial-temporal priors on large-scale datasets. In this paper, we aim to investigate whether such priors derived from a generative process are suitable for video recognition, and eventually joint optimization of generation and recognition. Building upon Stable Video Diffusion, we introduce GenRec, the first unified framework trained with a random-frame conditioning process so as to learn generalized spatial-temporal representations. The resulting framework can naturally supports generation and recognition, and more importantly is robust even when visual inputs contain limited information. Extensive experiments demonstrate the efficacy of GenRec for both recognition and generation. In particular, GenRec achieves competitive recognition performance, offering 75.8% and 87.2% accuracy on SSV2 and K400, respectively. GenRec also performs the best class-conditioned image-to-video generation results, achieving 46.5 and 49.3 FVD scores on SSV2 and EK-100 datasets. Furthermore, GenRec demonstrates extraordinary robustness in scenarios that only limited frames can be observed.
@@ 1 Introduction
While conceptually appealing, unifying video generation and recognition into a diffusion framework is non-trivial. Prior work either views the diffusion models as frozen feature extractors [9, 52, 39], or deconstructs them for new tasks while sacrificing their original generation capability [8]. One major challenge comes from their distinct training and inference processes. Diffusion models are typically optimized using corrupted inputs, optionally augmented with a single conditioning frame, to achieve unconditioned or image-conditioned generation during inference [27, 1]. In contrast, video recognition models require access to multiple frames to reason about temporal relationships and expect clean inputs during inference [48, 50]. Consequently, training a recognition model using corrupted videos and single-image conditions tends to suffer from inferior model optimization and a more significant training-inference gap.
@@ 3.3 Inference for Different Downstream Tasks
We can now replace the original score function with $p(\hat{\mathbf{z}}_{i})p(y|\hat{\mathbf{z}}_{i})$, in which $y$ denotes the conditional class, to get the conditional version of residual, denoted as $\epsilon^{*}_{\theta}(\hat{\mathbf{z}}_{i};\overline{\mathbf{\mathbf{z}}_{0}})$:
@@
$=\epsilon_{\theta}({\hat{\mathbf{z}}_{i}};\overline{\mathbf{\mathbf{z}}_{0}})-(t_{i-1}-\hat{t}_{i}){\hat{t}_{i}}{\nabla_{\hat{\mathbf{z}}_{i}}}\log p(y|\hat{\mathbf{z}}_{i})$ (16)
@@
$\epsilon^{*}_{\theta}(\hat{\mathbf{z}}_{i};\overline{\mathbf{\mathbf{z}}_{0}})=\epsilon_{\theta}({\hat{\mathbf{z}}_{i}};\overline{\mathbf{\mathbf{z}}_{0}})-\frac{(t_{i-1}-\hat{t}_{i}){\hat{t}_{i}}}{\sqrt{\hat{t_{i}}^{2}+\sigma_{data}}}{\nabla_{c(\hat{\mathbf{z}}_{i})}}\log p(y|c(\hat{\mathbf{z}}_{i}))$ (17)
@@ 5 Related Work
The great success of diffusion models in image generation has led to rapid advancements in video generation, including text-to-video generation [15, 2, 51], image&text-to-video generation [53, 18, 21], and video editing [4, 5, 26, 29, 12]. Many current works [51, 18] adapts the diffusion models from images to videos by incorporating temporal convolutions and attention mechanisms. One typical and excellent work, Stable Video Diffusion [1], follows the above description and has provided valuable foundations for generating high-quality, diverse, and temporally consistent videos. Different from the previous work, in our paper, we pursue not only the quality of generation, but also the unity of model generation capability and classification ability.
@@ A.1 Early Prediction on EK-100 and UCF-101
A.2 Case Study for Class-Conditioned Image-to-Video Generation and for Video Interpolation
@@
A.3 Limitations and Broader Impacts

### v2

@@ Abstract
Video diffusion models are able to generate high-quality videos by learning strong spatial-temporal priors on large-scale datasets. In this paper, we aim to investigate whether such priors derived from a generative process are suitable for video recognition, and eventually joint optimization of generation and recognition. Building upon Stable Video Diffusion, we introduce GenRec, the first unified framework trained with a random-frame conditioning process so as to learn generalized spatial-temporal representations. The resulting framework can naturally supports generation and recognition, and more importantly is robust even when visual inputs contain limited information. Extensive experiments demonstrate the efficacy of GenRec for both recognition and generation. In particular, GenRec achieves competitive recognition performance, offering 75.8% and 87.2% accuracy on SSV2 and K400, respectively. GenRec also performs the best on class-conditioned image-to-video generation, achieving 46.5 and 49.3 FVD scores on SSV2 and EK-100 datasets. Furthermore, GenRec demonstrates extraordinary robustness in scenarios that only limited frames can be observed. Code will be available at https://github.com/wengzejia1/GenRec.
Corresponding author.
@@ 1 Introduction
While conceptually appealing, unifying video generation and recognition into a diffusion framework is non-trivial. Prior work either views the diffusion models as frozen feature extractors [9, 55, 39], or deconstructs them for new tasks while sacrificing their original generation capability [8]. One major challenge comes from their distinct training and inference processes. Diffusion models are typically optimized using corrupted inputs, optionally augmented with a single conditioning frame, to achieve unconditioned or image-conditioned generation during inference [27, 1]. In contrast, video recognition models require access to multiple frames to reason about temporal relationships and expect clean inputs during inference [49, 51, 48]. Consequently, training a recognition model using corrupted videos and single-image conditions tends to suffer from inferior model optimization and a more significant training-inference gap.
@@ 3.3 Inference for Different Downstream Tasks
We substitute the score function ${\nabla_{\hat{\mathbf{z}}_{i}}}\log p_{\theta}(\hat{\mathbf{z}}_{i})$ with the conditional form ${\nabla_{\hat{\mathbf{z}}_{i}}}\log p_{\theta}(\hat{\mathbf{z}}_{i}|y)$, in which $y$ denotes the conditional class. By applying Bayes‘ Theorem, the original score function can be replaced by $p(\hat{\mathbf{z}}_{i})p(y|\hat{\mathbf{z}}_{i})$, and we can get the conditional version of residual, denoted as $\epsilon^{*}_{\theta}(\hat{\mathbf{z}}_{i};\overline{\mathbf{\mathbf{z}}_{0}})$:
@@
$=\epsilon_{\theta}({\hat{\mathbf{z}}_{i}};\overline{\mathbf{\mathbf{z}}_{0}})-(t_{i-1}-\hat{t}_{i}){\hat{t}_{i}}{\nabla_{\hat{\mathbf{z}}_{i}}}\log p(y|\hat{\mathbf{z}}_{i}).$ (16)
@@
$\epsilon^{*}_{\theta}(\hat{\mathbf{z}}_{i};\overline{\mathbf{\mathbf{z}}_{0}})=\epsilon_{\theta}({\hat{\mathbf{z}}_{i}};\overline{\mathbf{\mathbf{z}}_{0}})-\frac{(t_{i-1}-\hat{t}_{i}){\hat{t}_{i}}}{\sqrt{\hat{t_{i}}^{2}+\sigma_{data}}}{\nabla_{c(\hat{\mathbf{z}}_{i})}}\log p(y|c(\hat{\mathbf{z}}_{i})).$ (17)
@@ 4.2 Main Results
Table 6: SSV2 action recognition accuracy with different random seeds.
Seed 0 1 2 3 4 5 AVG
SSV2 Acc (%) 75.83 75.82 75.83 75.83 75.86 75.84 75.835 $\pm$ 0.0125
Noise incorporation during inference for video recognition.
In the inference stage for action recognition, GenRec applies a specific level of noise to video inputs before extracting visual features, as formulated in Equation 18. This added noise helps maintain consistency with the noisy training process. To further understand the influence of noise randomness on classification accuracy, we conducted an experiment on the SSV2 dataset, using multiple random seeds to generate the noise, as presented in Table 6. The results demonstrate remarkable consistency in accuracy across different random seeds, with a standard deviation of only 0.0125%. This minimal variation highlights the model’s robustness to noise fluctuations during inference, suggesting that the model’s performance remains stable despite noise introduced by random sampling. If fully deterministic outcomes are desired, fixing the random seed for noise sampling will eliminate any remaining variability and guarantee consistent predictions across runs.
@@ 5 Related Work
The great success of diffusion models in image generation has led to rapid advancements in video generation, including text-to-video generation [15, 2, 52, 54], image&text-to-video generation [56, 18, 21], and video editing [4, 5, 26, 29, 12, 53]. Many current works [52, 18] adapts the diffusion models from images to videos by incorporating temporal convolutions and attention mechanisms. One typical and excellent work, Stable Video Diffusion [1], follows the above description and has provided valuable foundations for generating high-quality, diverse, and temporally consistent videos. Different from the previous work, in our paper, we pursue not only the quality of generation, but also the unity of model generation capability and classification ability.
@@ 6 Conclusion
Acknowledgement
This work was supported in part by National Natural Science Foundation of China (#62032006). The authors would like to thank Rui Wang, Rong Bao, Rui Tian for their help and suggestions.
@@ A.1 Early Prediction on EK-100 and UCF-101
A.2 Training and Inference Time Cost Compared with Previous Classification Methods
We compare the training and inference time cost between MVD-H, Baseline II and GenRec, fairly on the same hardware resource: 4-nodes * 8 V100s, and same batch size. Baseline II refers to fully fine-tuning the original SVD model with classification supervision only.
As shown in the Table 9, GenRec and the Baseline II consumes more testing time than MVD-H. The difference primarily arises from the varying number of parameters. Our model is derived from a generative model, and the complexity of the generative task necessitates a larger number of parameters for effective learning. Compared with Baseline II, since GenRec requires additional decoder blocks for video generation training, the training time will get increased a little bit. As the additional up-sampling blocks will not be used when doing action recognition, GenRec shares the same testing time with Baseline II. It is worth noting that $\dagger$” in MVD-H is to highlight that MVD method would use repeated augmentation technique during training, but such augmentation will significantly increase the training time. Our approach does not need to use that augmentation. All the methods do the down-stream finetuning for 30 epochs.
Table 9: Comparison with recognition methods in terms of parameters, training time, and test time.
Method Trainable Total Training Training Test Time
Params Params Time Epochs (2 $\times$ 3 clips)
MVD-H 633M 633M 3038 s/Epoch 30 441 s
Baseline II 1.3B 1.9B 2954 s/Epoch 30 1500 s
GenRec 1.5B 2.1B 3751 s/Epoch 30 1500 s
A.3 Ablation on Loss Balance Ratio
We conduct an ablation study on the loss balance ratio $\lambda$ as shown in Equation 9. Results presented in Table 10 show that varying the loss ratio has a minor impact on action recognition accuracy. However, setting the ratio too low negatively affects the generation performance. In particular, when the ratio is set to zero, significant forgetting occurs, severely compromising the model’s generation ability.
Table 10: Ablation study on the effect of different loss balance ratios on SSv2 accuracy and FVD. Higher SSV2 accuracy and lower FVD indicate better performance.
Balance Ratio $\lambda$ 0 1 5 10 20
SSv2 Acc $\uparrow$ 75.6 75.5 75.6 75.8 75.5
SSv2 FVD $\downarrow$ 1579.2 52.0 47.4 46.5 46.9
A.4 Comparision with SVD Baseline
Comparing with the open-source Stable Video Diffusion directly is meaningful. We conduct frame prediction tests using SVD model on the SSV2 and Epic-Kitchen datasets, as shown in Table 11. Since the SVD model performs better at higher resolutions, we generated videos at 512x512 resolution and then downsampled them to 256x256 for FVD calculations. The results show that while SVD achieves competitive scores compared to previous state-of-the-art methods shown in Table 1, it is suboptimal compared to GenRec. This is likely due to SVD’s design for general scenarios. Baseline I represents our enhanced SVD baseline, which fine-tunes SVD on the target datasets for better results and fairer comparison with GenRec.
Table 11: Comparison of SSv2 FVD and Epic FVD scores in frame prediction tests.
Method SSv2 FVD Epic FVD
SVD 99.7 180.8
Baseline I 50.3 53.6
GenRec 46.5 49.3
A.5 Impact of Classification Training on Generative Ability
To demonstrate classification and generative can coexist well in our training without negatively impacting each other, we construct the comparison with "Baseline I (SVD baseline)", which adapts SVD on the target dataset without classification supervision, and then trains another classifier head freeze the generation backbone, ensuring no influence from classification supervision. We conduct comparisons on SSv2: FP (frame prediction) and CFP (class-condition frame prediction), as shown in Table 12. By comparing the FP results, the similar scores between the two methods conclude that classification loss does not degrade the model’s generative ability in our approach. Moreover, the CFP results indicate that our method, with its more accurate classification performance, can guide the model to achieve higher-quality video frame predictions.
Table 12: Comparison of SSv2 Acc and SSv2 FVD scores for Baseline I and GenRec methods in FP (frame prediction) and CFP (class-condition frame prediction) scenarios.
Method SSv2 Acc $\uparrow$ SSv2 FVD $\downarrow$
Baseline I (FP) - 55.5
GenRec (FP) - 55.3
Baseline I (CFP) 63.7 50.3
GenRec (CFP) 75.8 46.5
A.6 Case Study for Class-Conditioned Image-to-Video Generation and Video Interpolation
@@
A.7 Limitations and Broader Impacts